OpenStudy (anonymous):

Find the area of the region that is bounded by the given curve and lies in the specified sector

OpenStudy (anonymous):

\[r= \sqrt \sin \theta , 0\le \theta \le \frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

you know it is in the first quadrant and that you have to integrate

OpenStudy (anonymous):

i know that \[A= \int\limits_{0}^{\frac{ \pi }{ 2 }}\frac{ 1 }{ 2 }\sin \theta\] d theta

OpenStudy (anonymous):

do i use half angle identity?

OpenStudy (anonymous):

your integral is wrong

OpenStudy (anonymous):

it should be sin theta ^ (1/2)

OpenStudy (anonymous):

which would become integral of \[\frac{ 1 }{ 4 }\cos \theta\]

OpenStudy (anonymous):

well if you are treating r as if it where y then it would be the same thing as y = sqrt(sin theta)

OpenStudy (anonymous):

then you would convert it to sin theta ^ (1/2) and integrate from there

OpenStudy (anonymous):

\[A= \frac{ 1 }{ 2 }r^2 \rightarrow A= \frac{ 1 }{ 2 }(\sqrt{\sin \theta})^2d \theta \rightarrow A= \frac{ 1 }{ 2 }\sin \theta\]

OpenStudy (anonymous):

that is what i meant

OpenStudy (anonymous):

A= \[\int\limits_{0}^{\frac{ \pi }{ 2 }}\frac{ 1 }{ 2 }\sin \theta d \theta\]

OpenStudy (anonymous):

to my knowledge the way you find the area under a curve is by simply integrating, and assuming it is the area between two curves then you would subtract the top curve from the bottom curve when you integrate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then..... i'm not sure where to begin. I know the answer is 1/2

OpenStudy (anonymous):